# Axion Signal Search Using Hybrid Nuclear–Electronic Spin Systems

### v1

@@ Abstract
Figure 4: Projected $5\sigma$ sensitivity to axion-nucleon couplings. Shown are the calculated discovery thresholds on $g_{aNN}$ for a 209Bi donor ensemble under several standard magnetometry protocols: Ramsey (Blue), Hahn-echo (Orange), XY8 (Green), and continuous spin-lock ($T_{1\rho}$) detection (Black). All curves include realistic device parameters and noise models (Supplementary Tables II–III), and assume collective enhancement from an entangled ensemble of $N=10^{6}$ nuclear spins together with resonator gain $G_{\text{res}}(Q{=}10^{5})$. The effective signal field entering each protocol is $B_{e,\text{eff}}=F_{\text{ent}}G_{\text{res}}G_{\text{hyb}}B_{a}^{(N)}$. The purple shaded region represents the spin-lock Monte-Carlo distribution (median and 5%-95% band), which provides the best sensitivity across most of the mass range. The yellow band denotes the DFSZ model predictions, and the horizontal magenta region indicates the current SN1987A cooling bounds (95% C.L.). Part of the parameter windows are created from Ref. [45].
@@
Conclusion Hybrid nuclear-electronic upconversion enables axion wind sensitivity beyond the nuclear-inductive limit while preserving the required directional signatures. This approach is compatible with existing solid-state qubit platforms, enabling experimental validation without new fundamental infrastructure. By combining an axion-coupled nuclear ensemble with electronic readout, the architecture exploits strong axion-nucleus interactions, long nuclear coherence, and hyperfine-mediated signal transfer to a fast, high-fidelity electron channel. The framework links coherence times, hyperfine strength, and technical noise to achievable performance and identifies the sidereal modulation pattern as a robust detection fingerprint. Simulations based on silicon donor platforms show that bismuth, with its large hyperfine coupling, yields the greatest hybrid gain, and with collective enhancement and resonator coupling the scheme reaches projected $5\sigma$ DFSZ-level sensitivity across $m_{a}=10^{-16}$-$10^{-6}\,\text{eV}$ with Monte Carlo simulation. The approach is compatible with CMOS fabrication, scalable donor arrays, and parallel readout, and supports projected sensitivity down to $g_{aNN}\sim 10^{-25}$, positioning this hybrid solid-state platform as a competitive route toward next-generation dark-matter searches.

### v2

@@ Abstract
Figure 4: Projected $5\sigma$ sensitivity to axion–nucleon couplings ($g_{an}=mg_{aNN}$). Calculated discovery thresholds on $g_{aNN}$ for a 209Bi donor ensemble using standard magnetometry protocols: Ramsey (Blue), Hahn-echo (Orange), XY8 (Green), and continuous spin-lock (Black). Curves include realistic device parameters and noise models [1], and assume collective enhancement from an entangled ensemble of $N=10^{6}$ nuclear spins plus resonator gain $G_{\mathrm{res}}(Q=10^{5})$. The effective signal field is $B_{e,\mathrm{eff}}=F_{\mathrm{ent}}G_{\mathrm{res}}G_{\mathrm{hyb}}B_{a}^{(N)}$, where $F_{\mathrm{ent}}$ parameterize quantum ensemble enhancement, e.g., $F_{\mathrm{ent}}=\sqrt{N}$ for uncorrelated spins at the SQL; $F_{\mathrm{ent}}=N$ for an ideal fully entangled state. Purple shading: Best spin-lock Monte-Carlo (95% C.L.) with the noise model in SM Tab.IV, giving best sensitivity over most masses. Yellow band: DFSZ model predictions; green Neutron star cooling region: current SN1987A cooling lower bounds. Portions of the parameter windows are from Ref. [46].
@@
Conclusion- In summary, we propose a hybrid nuclear–electronic upconversion scheme that exceeds the nuclear-inductive limit for axion-wind sensing while retaining directional (sidereal) signatures. Using a long-lived, axion-coupled nuclear ensemble and hyperfine transfer to a fast electronic readout, the method is compatible with silicon-donor solid-state qubit platforms and requires no new fundamental infrastructure. Monte Carlo simulations indicate the largest gain for bismuth donors, and with collective enhancement and resonator coupling we project $5\sigma$ DFSZ-level sensitivity over $m_{a}=10^{-16}-10^{-6}$ eV, reaching $g_{aNN}\sim 10^{-25}$.